OpenStudy (anonymous):

In Angle ABC, C is a right angle. Find the remaining sides and angles. Round your answers to the nearest tenth. a = 3, c = 19 A. M

OpenStudy (anonymous):

Im stuck between a and b

OpenStudy (anonymous):

(c^2 - a^2 = b^2) a=3, c=19 i plugged in, so 3^2=9, 19^2 =361 the in subtracted, so i got 352. now I'm kind of stuck on how to get 18.8

OpenStudy (campbell_st):

draw a diagram  find b using pythagoras theorem \[19^2 = 3^2 + b^2 ...or..... b^2 = 19^2 - 3^2\] then B can be found using any trig ratio... and A can be found using angle sum of a triangle hope it helps

OpenStudy (campbell_st):

take the square roots of 352

OpenStudy (anonymous):

so 18.7. and now i round it? @campbell_st

OpenStudy (campbell_st):

well 18.76... rounds to 18.8

OpenStudy (campbell_st):

 now pick an angle and trig ratio

OpenStudy (anonymous):

okay i think i got it. i think I'm just going to go with a. but thank you!

OpenStudy (anonymous):

@campbell_st hes the math KING

OpenStudy (campbell_st):

I think A is a good choice as the triangle is almost isosceles... 1 side 19 and 1 18.8